Skip to content
dronacharya.ai
← JEE Main 2017 · 8 Apr · Shift 1
MathematicsSingle correct+4 / -1Question 86 of 90

JEE Main 2017 · 8 Apr · Shift 1, Mathematics question 86

Three persons P, Q and R independently try to hit a target. If the probabilities of their hitting the target are 34\dfrac34, 12\dfrac12 and 58\dfrac58 respectively, then the probability that the target is hit by P or Q but not by R is :

Choose one option
Answer and worked solution

Official answer: A

P(P or Q)=1−14⋅12=78P(\text{P or Q})=1-\frac14\cdot\frac12=\frac78 and P(not R)=38P(\text{not R})=\frac38, so the probability is 2164\frac{21}{64}.